OpenStudy (anonymous):

(4x^2-y)^3 does that equal 64^8-15x^4-8x^2y^2-y?

OpenStudy (goformit100):

use binomial expansion or pascal trianle metod

OpenStudy (anonymous):

i did that. is that the right answer? or what should it be?

OpenStudy (goformit100):

Yes it will be the correc t answer

OpenStudy (anonymous):

Simply use formula.. (a-b)^3 = a^3 - 3(a^2)b - 3 a(b^2) - b^3 put 4x^2 = a and y= b in above formula..

OpenStudy (anonymous):

i did that! i just want to know what i did wrong in my answer if i got it wrong?

OpenStudy (anonymous):

ok... (4x^2)^3 - 3(4x^2)^2 (y) - 3 (4x^2)(y)^2 - (y)^3 = 64x^6 - 3(16x^4)(y) - 3(4x^2)(y^2) - y3 = 64x^6 - 48x^4 y - 12 x^2 y^2 - y^3

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):

got it now..??

OpenStudy (anonymous):

yes! i noticed what i did wrong. could you help me with another problem please?

OpenStudy (anonymous):

please post that question and message me its link..

OpenStudy (anonymous):

ok